OpenStudy (anonymous):

what volume of water must be added to 500cm3 of a 0.12M solution of sodium carbonate to give a 0.040M solution?

OpenStudy (colouringme):

\[\frac{ n }{ V } = M\] this is the equation you need to know. you know the current M and V. so what's the current n?

OpenStudy (colouringme):

V in liter

OpenStudy (anonymous):

M1V1=M2V2 500 X 0.12 = 0.04 X V2 500 X 0.12 / 0.04 = V2 1500 = V2 now total volume should be 1500 cm3 is required. you have already 500 cm3 as a parent solution, (1500-500=1000) so u have to add 1000 cm3 in 500 cm3 solution of 0.12 M

OpenStudy (colouringme):

yeah, I forgot that equation exist, hahah..

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

thanks!!